The average price of 3 mugs is $4, one of the mugs cost $p and another mug cost $3...express the price of the 3rd mug in terms of p...: The average price of 3 mugs is $4, one of the mugs cost $p and another mug cost $3...express the price of the 3rd mug in terms of p...
Answer by jojo14344(1029)

By Averaging:
(M[1]+M[2]+M[3])/3=4 ------> working eqn
Price:
M[1] = $p1
M[2] = $3
M[3] = $p2
Substituting:
(p[1]+3+p[2])/3=4
p[1]+3+p[2]=3*4
p[2]=12-3-p[1]
p[2]=9-p[1] -----------------> ANSWER, price of 3rd Mug
. By knowing price$ of M[1], you can get price$ of M[3] by subtacting from $9.00
Thank you,
Jojo